06 March 2013

Ridge Lab team defeats snow drifts

Eureka continued to enjoy temperatures near -42°C today. Up at the PEARL Ridge Lab, relatively mild temperatures continued, at -22°C. The intense winds of yesterday were gone, though the snow drifts along the road remained.

The Ridge Lab team shoveled their way through the snow drifts and continued with their measurements. While we now enjoy full workdays of sunlight, cloudy skies prevented solar measurements in mid-morning, early-afternoon, and late afternoon. In between cloudy conditions, Dan and Joseph took 46 measurements with the Bruker and Debora took 25 measurements with PARIS. Debora also continued tests with the new WACCM a prioris for several gases.

Over in the UV-Vis Lab, SAOZ, PEARL-GBS, and UT-GBS are working well. Xiaoyi soldered extension cables for the GBS' new temperature sensors, and set up a new temperature monitor. He also investigated how to set up a cooling system for the PEARL-GBS. The existing cooling system no longer fits because of PEARL-GBS' newly installed wedged CCD mount. Xiaoyi also sent a brief report to Dr. Goutail and Dr. Walker to discuss the future of SAOZ at PEARL.

The Ridge Lab team applied their nascent filmmaking skills this afternoon, recording segments of video for an upcoming CANDAC outreach project.

Volodya came up to the Ridge Lab this morning to check MAESTRO and SPS one last time before he leaves Eureka. He was surprised to see very dim spectra. Snow had built up inside the viewing ports, undoubtedly a result of yesterday's windy conditions. The last week was so calm, and there had been so little snow on the ground, that he never needed to clear the ports. This had been a daily routine in previous years. Once Jim cleared the ports with compressed air the spectra returned to normal. Meanwhile, Volodya completed the operator's instructions and the instructions for packing the instruments at the end of the campaign. He also contacted Yan to arrange access to the MAESTRO and SPS data from Toronto. Tom and Volodya will continue to monitor the instruments' health remotely. Jim registered the MAESTRO PC so that it has internet access for emergency situations. Volodya wishes the remaining team members a good two weeks of observations and a great time in Eureka.

Inside and around 0PAL, Zen and Emily also recorded outreach videos discussing their instruments and work. Chris ran the CRL all day and all night again. Emily was pleased that a cloud passed overhead; she will use this event to analyze her calibrations and make adjustments as needed. Once it arrived, Zen helped unload the plane and unpack supplies. Later in the day, he also packed for his trip back home tomorrow.

Today's ACE ozonesonde flight occurred at 18:15 (local time) and reached an altitude of 31,197 m (7.3 hPa).

Tonight the team gathered in the recreation room to bid farewell to Volodya and Zen, who depart in the morning. Also leaving are CANDAC Principal Investigator, Jim Drummond, and a few of our Environment Canada friends.

Dan

Back to Daily Reports
Back to Home